OpenStudy (vishal_kothari):

Each of the 11 letters A, H, I, M, O, T, U, V, W, X and Z appears same when looked at in a mirror. They are called symmetric letters. Other letters in the alphabet are asymmetric letters. How many three letter computer passwords can be formed (no repetition allowed) with at least one symmetric letter?

OpenStudy (anonymous):

There are 26-11 = 15 asymmetric letters; Your answer: \(26 \times 25 \times 24 - 15 \times 14 \times 13 \)